A pair of dice is cast and the numbers that show on the dice are observed. Find the probability that one die shows 6 and the other is a number less than 3.



1
Expert's answer
2021-11-29T05:05:51-0500

Since each die has 6 faces, the probability of getting 6 points on the first die is

"{p_1} = \frac{1}{6}"

The probability of getting a number less than 6 is

"{p_2} = \frac{2}{6} = \frac{1}{3}" (1 or two points).

Then the wanted probability is

"p = {p_1}{p_2} = \frac{1}{6} \cdot \frac{1}{3} = \frac{1}{{18}}"

Answer: "p = \frac{1}{{18}}"